OpenStudy (firejay5):

Solve the absolute value and show work please and there should be 2 answers: -12 |9x + 1| = 144

OpenStudy (anonymous):

To solve this, first divide both sides by -12 isolate the absolute value.

OpenStudy (anonymous):

So |9x + 1| = -12

OpenStudy (anonymous):

This is not possible Absolute values can never be negative. are you sure you wrote the problem correctly?

OpenStudy (firejay5):

yes

OpenStudy (firejay5):

the answer could be no solution as well if it doesn't work

OpenStudy (anonymous):

yeah no solution

OpenStudy (firejay5):

I think this one is no solution as well: |A - 3| -14 = -6

OpenStudy (anonymous):

no this has a solution |A - 3| = 8

OpenStudy (firejay5):

is there 1 or 2 answers for it

OpenStudy (anonymous):

2

OpenStudy (anonymous):

to solve these, you think of it like A-3 = 8 or A-3 = -8

OpenStudy (firejay5):

that's actually what we learned in class today that's funny

OpenStudy (anonymous):

well that's probably why you have to solve inequalities. but do you understand how to solve it now?

OpenStudy (firejay5):

Yea it confuses me a little, but I got it now